OpenStudy (anonymous):

Help? Step by step.. Give the slope and the y intercept of the line -2x-4y=5.

OpenStudy (anonymous):

rearrange into y = ax + b -2x - 4y = 5 - 4y = 2x + 5 y = -2/4x -5/4 y = -1/2x - 5/4 the gradient/slope is the coefficient of x and the y-intercept can be calculated when x = 0 or u can see it is the last term. y = -1/2(0) - 5/4 y = -5/4